OpenStudy (anonymous):

whats 1 8/9 as a percent

jimthompson5910 (jim_thompson5910):

1 8/9 = 1+8/9 = 1+0.8888 = 1.8888 = 188.88%

OpenStudy (anonymous):

Convert into a decimal, multiply by 100, add a percent sign.